OpenStudy (anonymous):

If y=(x^3)+5x and dx/dt=7. Find dy/dt when x=3

OpenStudy (jamesj):

By the chain rule \[ \frac{dy}{dt} = \frac{dy}{dx}\frac{dx}{dt} \] See what to do now?

OpenStudy (anonymous):

What am I supposed to plug in?

OpenStudy (jamesj):

Well, you are given dx/dt, so that's one piece. And you can calculate for yourself dy/dx, no?

OpenStudy (phi):

y= x^3+5x Can you find dy/dx ? (use the power rule) Once you have dy/dx, evaluate it at x=3 now put all the pieces together (see JamesJ above)